# Do LLMs Encode Functional Importance of Reasoning Tokens ?

### v1

@@ Abstract
Large language models solve complex tasks by generating long reasoning chains, achieving higher accuracy at the cost of increased computational cost and reduced ability to isolate functionally relevant reasoning. Prior work on compact reasoning shortens such chains through probabilistic sampling, heuristics, or supervision from frontier models, but offers limited insight into whether models internally encode token-level functional importance for answer generation. We address this gap diagnostically and propose greedy pruning, a likelihood-preserving deletion procedure that iteratively removes reasoning tokens whose removal minimally degrades model likelihood under a specified objective, yielding length-controlled reasoning chains. We evaluate pruned reasoning in a distillation framework and show that students trained on pruned chains outperform a frontier-model–supervised compression baseline at matched reasoning lengths. Finally, our analysis reveals systematic pruning patterns and shows that attention scores can predict greedy pruning ranks, further suggesting that models encode a nontrivial functional importance structure over reasoning tokens. [^1] Code: github.com/greedy-token-pruner/greedy-token-pruner; Demo: greedy-token-pruner.github.io.
@@ 4.2 Results
Across all settings, greedy pruning yields the strongest student performance among pruning-based methods at matched keep fractions. Accuracy improves smoothly as the keep fraction increases, indicating graceful degradation under compression. On GSM8K, students trained on greedy-pruned reasoning outperform TokenSkip, H2O, Surprisal, and Uniform across all keep fractions, for both Llama3.1-8B and Qwen2.5-7B teacher–pruner pairs (Figures 2(a), 2(b)). Uniform and Surprisal drop sharply at aggressive pruning, while TokenSkip improves with higher keep fractions but remains below greedy pruning. Zero-shot student performance is shown for reference and is consistently worse than all pruning-based distillation methods. Similar trends hold on the more challenging MMLU-Pro and MATH benchmarks (Figures 2(c), 2(d)): despite lower absolute accuracy, greedy pruning remains best at matched reasoning lengths. Additional results with Mistral-7B as the student show the similar trends (Appendix C).
Overall, pruned reasoning supports effective distillation across datasets, models, and keep fractions, outperforming the frontier-model–supervised TokenSkip baseline. This suggests that greedy pruning preserves tokens important for answer generation and motivates a closer analysis of the induced importance structure, which we study next.
@@ 5.2 Dynamics of Pruning Ranks
Dynamic local rankings consistently outperform both a Frozen=1.0 baseline, which reuses rankings computed at $\rho=1.0$ to simulate a fixed global ordering, and a random baseline. The widening gap between dynamic and frozen rankings, especially at intermediate keep fractions, shows that greedy pruning does not expose a static importance ordering early in pruning. Instead, token importance is continually re-evaluated as reasoning chain contracts, with later-stage local rankings remaining most predictive of imminent pruning decisions. This dynamic behavior highlights greedy pruning as an adaptive process rather than one governed by a fixed global notion of importance.
@@ 5.3 Predicting Pruning Ranks from Attention
The surrogate is trained to predict the post-deletion likelihood $L^{del}$ associated with removing each token under greedy pruning, using a Pearson correlation objective that emphasizes relative importance rather than absolute scale. This setup allows us to test whether attention patterns encode sufficient information to predict the pruning order of reasoning tokens. Using only 200 GSM8K examples for training with Llama3.1-8B as the teacher, the surrogate achieves a Pearson correlation of 0.88 with true post-deletion likelihoods on a held-out set of 1,000 GSM8K examples. These results indicate that attention patterns encode information strongly predictive of greedy pruning decisions, and suggest a promising direction for incorporating such signals into more efficient pruning or search procedures.
@@ 6 Discussion & Future Directions
Greedy pruning can be viewed as a structured deletion operator for compressing reasoning chains under likelihood-based objectives, enabling controlled exploration of reasoning space in LLMs. Under this view, greedy pruning represents one member of a broader family of reasoning-space operators, which could be extended beyond deletion to token addition or replacement under alternative objectives, enabling finer-grained control over reasoning trajectories. Such operators offer a structured means of exploring reasoning space rather than relying on temperature-based sampling.
Beyond compression, greedy pruning serves as a diagnostic probe of model-internal reasoning, revealing which parts of a model’s own generated reasoning are functionally necessary for producing correct answers. Our analyses further show that pruning behavior depends on both the pruning objective and pruner strength, suggesting that pruning structure itself may reflect model capability and internal organization, and represents an interesting direction for further study.
In addition, pruning ranks are strongly predictable from attention patterns, indicating that token importance signals may be accessible during generation. This observation points toward inference-time pruning mechanisms, as explored by recent work Yang et al. (2025); Monea et al. (2025), which discard completed reasoning segments to manage context length. While we focus on likelihood-preserving post-hoc pruning as a diagnostic tool, extending similar objectives to inference-time pruning within memory-efficient reasoning systems, or integrating pruning into training-time curricula for distillation, remains an open direction.
@@ 7 Conclusion
We investigate whether LLMs encode token-level functional importance for answer generation within their reasoning chains. Using greedy pruning, a likelihood-preserving deletion procedure, we show that models can systematically compress reasoning while retaining information critical for answer generation, and that students trained on such pruned reasoning outperform multiple baselines at matched reasoning lengths. Our analyses reveal a stable and interpretable functional structure under pruning, with symbolic computation preferentially preserved and supporting linguistic scaffolding pruned earlier, and further show that pruning behavior depends on the pruning objective and pruner strength. We also demonstrate that pruning ranks are dynamically reshaped as context contracts and are strongly predictable from attention patterns, indicating that models encode internal signals of reasoning-token importance. Together, these findings establish greedy pruning as a principled diagnostic method for exposing token-level importance structure in model-generated reasoning.
@@
We use likelihood-based greedy pruning to expose token-level importance structure among reasoning tokens encoded in LLMs. However, greedy pruning is computationally expensive, as it requires repeated likelihood evaluations for candidate token deletions, which can limit scalability to larger models or longer reasoning chains. Several strategies could mitigate this cost. First, prefix caching [^2] https://docs.vllm.ai/en/stable/design/prefix_caching/ can reuse KV caches for shared prefixes across likelihood evaluations, since multiple deletion candidates share long prefixes. Second, induced pruning ranks could enable deleting multiple tokens per step instead of recomputing deletion scores at every iteration, as motivated by the strong alignment between local pruning ranks and subsequent pruning decisions observed in Section 5.2. Third, attention-based or learned surrogate models could approximate importance scores without full likelihood recomputation, as suggested by the strong correlation between post-deletion likelihoods and attention-based features observed in Section 5.3.
Beyond efficiency, at extreme prune ratios, pruned reasoning traces may become difficult to interpret even when task accuracy is largely preserved. Our work does not address how to determine the maximum compression level at which reasoning remains comprehensible or suitable for human inspection, and developing principled criteria for reasoning interpretability under aggressive compression remains an open problem.
Finally, our experiments are restricted to relatively smaller instruction-tuned LLMs and do not extend to DeepSeek-style Guo et al. (2025b) large reasoning models due to computational constraints. However, we find that weaker pruners induce qualitatively poorer pruning ranks (Section 5.1), suggesting that stronger models may capture more reliable token-importance signals. As a result, the qualitative trends identified in this work may strengthen rather than weaken at larger model scales.
@@ B.1 Datasets
  • GSM8K. The original dataset provides 7.47K training examples and 1.32K test examples. We randomly partition the training split into 6.97K training examples and 0.5K development examples, while keeping the original test split intact.
  • MMLU-Pro. The original dataset does not provide a standard training split. We therefore randomly partition the available 12K examples into 10K training, 0.8K development, and 1.23K test examples.
  • MATH. The dataset provides 12K training examples and a 500-example test set. We randomly split the training data into 11.2K training examples and 0.8K development examples, and evaluate on the original MATH-500 test set.
@@ B.4 Training and Inference
  • Distillation Training (Student). All supervised fine-tuning (SFT) experiments are implemented using the Axolotl framework Axolotl maintainers and contributors (2023). We sweep learning rates in $\{1\times 10^{-6},\,3\times 10^{-6},\,5\times 10^{-6},\,1\times 10^{-5},\,3\times 10^{-5},\,5\times 10^{-5}\}$ using a cosine decay scheduler. Across all Llama2-7B training runs, a learning rate of $3\times 10^{-5}$ consistently yields the best performance, whereas Mistral-7B prefers a lower learning rate of $5\times 10^{-6}$. All student models are trained with an effective batch size of 28 for 1,200 gradient steps.

### v3

@@ Abstract
Large language models solve complex tasks by generating long reasoning chains, achieving higher accuracy at the cost of increased computational cost and reduced ability to isolate functionally relevant reasoning. Prior work on compact reasoning shortens such chains through probabilistic sampling, heuristics, or supervision from frontier models, but offers limited insight into whether models internally encode token-level functional importance for answer generation. We address this gap diagnostically and propose greedy pruning, a likelihood-preserving deletion procedure that iteratively removes reasoning tokens whose removal minimally degrades model likelihood under a specified objective, yielding length-controlled reasoning chains. We evaluate pruned reasoning in a distillation framework and show that students trained on pruned chains outperform a frontier-model–supervised compression baseline at matched reasoning lengths. Finally, our analysis reveals systematic pruning patterns and shows that attention scores can predict greedy pruning ranks, further suggesting that models encode a nontrivial functional importance structure over reasoning tokens. [^1] Code: github.com/iamjanvijay/greedy-token-pruner; Demo: iamjanvijay.github.io/greedy-token-pruner.
@@ 4.2 Results
Across all settings, greedy pruning yields the strongest student performance among pruning-based methods at matched keep fractions. Accuracy improves smoothly as the keep fraction increases, indicating graceful degradation under compression. On GSM8K, students trained on greedy-pruned reasoning outperform TokenSkip, H2O, Surprisal, and Uniform across all keep fractions, for both Llama3.1-8B and Qwen2.5-7B teacher–pruner pairs (Figures 2(a), 2(b)). Uniform and Surprisal drop sharply at aggressive pruning, while TokenSkip improves with higher keep fractions but remains below greedy pruning. Zero-shot student performance is shown for reference and is consistently worse than all pruning-based distillation methods. Similar trends hold on the more challenging MMLU-Pro and MATH benchmarks (Figures 2(c), 2(d)): despite lower absolute accuracy, greedy pruning remains best at matched reasoning lengths. Additional results with Mistral-7B as the student show similar trends (Appendix C). Overall, pruned reasoning supports effective distillation across datasets, models, and keep fractions, consistently outperforming the frontier-model–supervised TokenSkip baseline.
We further analyze three key aspects of our compressed-reasoning distillation approach. First, we quantify the computational cost of greedy pruning (Appendix B.5), showing that while the naive implementation scales superlinearly with sequence length, it is a one-time offline preprocessing step and admits practical optimizations. Second, we evaluate the readability and structural integrity of pruned traces (Appendix E), finding that even at 30% token reduction, degradation is primarily surface-level, with semantic coherence and mathematical state largely preserved. Third, we study the behavior of students trained on pruned data (Appendix F), observing that they learn to generate proportionally shorter reasoning traces without meaningful loss in semantics or structure.
Together, these results indicate that greedy pruning removes redundant tokens while preserving those critical for answer generation, motivating a closer analysis of the induced importance structure, which we study next.
@@ 5.2 Dynamics of Pruning Ranks
Dynamic local rankings consistently outperform both a Frozen=1.0 baseline, which reuses rankings computed at $\rho=1.0$ to simulate a fixed global ordering, and a random baseline. The widening gap between dynamic and frozen rankings, especially at intermediate keep fractions, shows that greedy pruning does not expose a static importance ordering early in pruning. Instead, token importance is continually re-evaluated as the reasoning chain contracts, with later-stage local rankings remaining most predictive of imminent pruning decisions. This dynamic behavior highlights greedy pruning as an adaptive process rather than one governed by a fixed global notion of importance.
@@ 5.3 Predicting Pruning Ranks from Attention
The surrogate is trained to predict the post-deletion likelihood $L^{del}$ associated with removing each token under greedy pruning, using a Pearson correlation objective that emphasizes relative importance rather than absolute scale. This setup allows us to test whether attention patterns encode sufficient information to predict the pruning order of reasoning tokens. Using only 200 GSM8K examples for training with Llama3.1-8B as the teacher, the surrogate achieves a Pearson correlation of 0.88 with true post-deletion likelihoods on a held-out set of 1,000 GSM8K examples. These results indicate that attention patterns encode information strongly predictive of greedy pruning decisions and suggest a promising direction for incorporating such signals into more efficient pruning or search procedures.
@@ 6 Discussion & Future Directions
Beyond compression, greedy pruning also serves as a diagnostic probe of model-internal reasoning. By selectively removing tokens while preserving output likelihood, it reveals which parts of a model’s generated reasoning are functionally necessary for producing correct answers. Our results (Section˜5.1) further show that pruning behavior depends on both the pruning objective and the pruner’s capacity, suggesting that the induced importance structure reflects the properties of the underlying model. Understanding how this structure varies across models, scales, and training procedures opens a promising avenue for future work.
Beyond these diagnostic implications, greedy pruning can be viewed as a structured deletion operator for compressing reasoning chains under likelihood-based objectives, offering a controlled mechanism for exploring reasoning space in LLMs. From this perspective, it represents one instance of a broader class of reasoning-space operators. Extending this framework beyond deletion to operations such as token insertion or replacement under alternative objectives could enable finer-grained manipulation of reasoning trajectories, providing a more principled alternative to purely sampling-based control.
The tractability of such extensions is further supported by our finding that pruning ranks are strongly predictable from attention-based signals, suggesting that token-level importance may be partially accessible during generation. This connects naturally to recent work on inference-time pruning Yang et al. (2025); Monea et al. (2025), which discards completed reasoning segments to manage context length. While our work focuses on likelihood-preserving post-hoc pruning as an analysis and distillation tool, extending similar objectives to inference-time settings or incorporating pruning signals into training-time curricula represents a natural next step toward more efficient and controllable reasoning systems.
Together, these directions suggest that pruning-based objectives may serve as a unifying interface for analyzing and shaping reasoning in LLMs, not merely as a compression tool but as a principled lens into model cognition.
@@ 7 Conclusion
We investigate whether LLMs encode token-level functional importance for answer generation within their reasoning chains. Using greedy pruning, a likelihood-preserving deletion procedure, we show that models can systematically compress reasoning while retaining information critical for answer generation, and that students trained on such pruned reasoning outperform multiple baselines at matched reasoning lengths. Our analyses reveal a stable and interpretable functional structure under pruning, with symbolic computation preferentially preserved and supporting linguistic scaffolding pruned earlier. Furthermore, we show that pruning behavior depends on the pruning objective and pruner strength. We also demonstrate that pruning ranks are dynamically reshaped as context contracts and are strongly predictable from attention patterns, indicating that models encode internal signals of reasoning-token importance. Together, these findings establish greedy pruning as a diagnostic method for exposing token-level importance structure in model-generated reasoning.
@@
We study likelihood-based greedy pruning as a tool for exposing token-level importance structure in LLM reasoning. A primary limitation is computational cost: greedy pruning requires repeated likelihood evaluations over candidate token deletions, leading to superlinear scaling in sequence length and making naive implementations expensive for long reasoning chains. In our setting, this cost is incurred as a one-time offline preprocessing step, amortized over downstream student training. Importantly, our goal is not to provide a universally scalable pruning algorithm but to analyze token-level importance at realistic reasoning lengths (e.g., $\leq$500 tokens), where the method remains empirically tractable.
Several directions could reduce this cost. First, prefix caching [^2] https://docs.vllm.ai/en/stable/design/prefix_caching/ can reuse KV states for shared prefixes across deletion candidates, yielding constant-factor speedups. Second, the stability of local pruning ranks (Section˜5.2) suggests that multiple tokens could be removed per iteration, reducing the number of recomputation steps. Third, attention-based or learned surrogate models could approximate token importance without full likelihood recomputation, supported by the strong correlation between post-deletion likelihoods and attention-derived features (Section˜5.3). A detailed analysis of computational costs and scaling behavior is provided in Section˜B.5.
Beyond efficiency, aggressive pruning can degrade the readability of reasoning traces. While moderate pruning primarily introduces formatting-level artifacts rather than loss of semantic or mathematical structure (Appendix˜E), we do not characterize the precise threshold at which reasoning becomes unsuitable for human interpretation. Developing principled criteria for reasoning interpretability under compression remains an open problem.
Our analysis is further restricted to teacher-generated, correctness-filtered reasoning traces. How pruning behaves on incorrect or low-quality reasoning, where likelihood signals may be less aligned with functional importance, remains an open question.
Our experiments are conducted on moderately sized instruction-tuned LLMs and do not extend to very large reasoning-focused models (e.g., DeepSeek-style Guo et al. (2025b)) due to computational constraints. That said, our ablations indicate that pruning quality is strongly governed by pruner strength (Section˜5.1), suggesting that larger models may yield more reliable token-importance estimates and that the qualitative trends observed here are likely to persist or strengthen at scale.
Acknowledgments
This work used the DeltaAI system at the National Center for Supercomputing Applications [award OAC 2320345] through allocation CIS251207 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ B.1 Datasets
  • GSM8K. The original dataset provides 7.47K training examples and 1.32K test examples. We randomly partition the training split into 6.97K training examples and 0.5K development examples while keeping the original test split intact. The resulting split is released at iamjanvijay/gsm8k.
  • MMLU-Pro. The original dataset does not provide a standard training split. We therefore randomly partition the available 12K examples into 10K training, 0.8K development, and 1.23K test examples. The resulting split is released at iamjanvijay/MMLU-Pro.
  • MATH. The dataset provides 12K training examples and a 500-example test set. We randomly split the training data into 11.2K training examples and 0.8K development examples and evaluate on the original MATH-500 test set. The resulting split is released at iamjanvijay/openaimath.
@@ B.4 Training and Inference
  • Distillation Training (Student). All supervised fine-tuning (SFT) experiments are implemented using the Axolotl framework Axolotl maintainers and contributors (2023). We sweep learning rates in $\{1\times 10^{-6},\,3\times 10^{-6},\,5\times 10^{-6},\,1\times 10^{-5},\,3\times 10^{-5},\,5\times 10^{-5}\}$ using a cosine decay scheduler. Across all Llama2-7B training runs, a learning rate of $3\times 10^{-5}$ consistently yields the best performance, whereas Mistral-7B prefers a lower learning rate of $5\times 10^{-6}$. All student models are trained with an effective batch size of 28 for 1,200 gradient steps, and the best checkpoint is selected using validation set performance.
B.5 Computational Cost and Scaling of Greedy Pruning
Greedy pruning incurs a nontrivial computational cost due to repeated likelihood evaluations under token deletions. For a reasoning trace with $n$ tokens and a target keep fraction of $\rho$, the algorithm evaluates $\mathcal{O}((1-\rho)n)$ pruning steps. At each step, all remaining tokens are considered deletion candidates, resulting in a total of approximately $\mathcal{O}((1-\rho)n^{2})$ forward passes. Each candidate evaluation requires scoring the full sequence under the pruner model. Under standard transformer implementations, this corresponds to $\mathcal{O}(n^{2})$ self-attention costs per forward pass. Thus, the naive worst-case FLOP complexity is $\mathcal{O}((1-\rho)n^{4})$, which can become expensive for longer reasoning traces.
[…]
Effect of sequence length.
The quadratic dependence on sequence length implies that greedy pruning becomes increasingly expensive for longer generations (e.g., >1>1K tokens). While our experiments demonstrate feasibility for moderately long reasoning traces (up to $\sim$500 tokens), scaling to substantially longer outputs would require additional approximations.
Optimization opportunities.
As previously mentioned, several optimizations can reduce the practical runtime:
  • Prefix KV-cache reuse. For deletion candidates that share common prefixes, key–value states can be reused exactly, reducing redundant computation. While this does not change asymptotic complexity, it yields a meaningful constant-factor speedup.
  • Block or local multi-token deletion. Instead of removing a single token per step, one can remove a small set of high-ranked tokens and recompute importance locally. This reduces the number of pruning iterations at the cost of introducing approximation.
  • Surrogate importance models. Section˜5.3 shows that attention-based features correlate strongly with pruning ranks. Lightweight surrogate models can approximate token importance and reduce the number of exact likelihood evaluations.
Discussion.
We emphasize that greedy pruning is designed as an offline analysis and data construction procedure, rather than a test-time algorithm. Our goal is to expose token-level functional structure and enable efficient distillation, rather than to provide a universally scalable pruning method for arbitrarily long generations. Improving scalability remains an important direction for future work.
@@ D.3 Additional Results
Appendix E Comprehensibility and SFT Suitability of Pruned Reasoning
We evaluate the structural and semantic quality of reasoning traces produced by greedy pruning. While our method optimizes a likelihood-based objective, it is important to verify that the resulting supervision remains suitable for training and interpretable to humans.
E.1 Evaluation Protocol
We evaluate pruned reasoning traces on GSM8K using Llama3.1-8B as the teacher and pruner. For each keep fraction $\rho\in\{0.7,0.8,0.9\}$, we sample 1,000 reasoning traces after pruning.
Each trace is evaluated along five dimensions:
  • Corruption: surface-level degradation of text.
  • Structural integrity: preservation of a stepwise reasoning structure.
  • Semantic coherence: logical consistency and readability.
  • Mathematical state preservation: whether key variables, intermediate quantities, and final answers are present and consistent.
  • Suitability for SFT: whether the trace is usable as supervision for fine-tuning.
Annotations are generated using an LLM-based evaluator (temperature =0=0) under a fixed JSON schema to ensure deterministic outputs. To validate annotation quality, we manually inspect 50 randomly sampled traces at $\rho=0.7$.
E.2 Annotation Rubric
Each dimension is labeled using a discrete scale:
Corruption
  • OK: fully readable text.
  • Minor: surface artifacts such as missing punctuation, dropped function words, or formatting noise; meaning remains recoverable.
  • Severe: fragmented or malformed text that is difficult to parse.
Structural Integrity
  • Intact: step structure preserved.
  • Mildly broken: minor step merging or header loss; order remains recoverable.
  • Broken: reasoning flow is difficult to reconstruct.
Semantic Coherence
  • OK: logically understandable.
  • Minor: connective or phrasing gaps, but reasoning remains clear.
  • Severe: logical flow is disrupted.
Mathematical State Preservation
  • Yes: key quantities and equations are present and numerically consistent.
  • Borderline: minor omissions such as units; state remains inferable.
  • No: critical quantities are missing or inconsistent.
Suitability for SFT
  • Yes: directly usable supervision.
  • Borderline: usable with minor cleanup.
  • No: too degraded for training.
E.3 Results
Table 3 reports results across keep fractions.
Metric Category $\rho=0.7$ $\rho=0.8$ $\rho=0.9$ Corruption OK 3.9% 5.1% 22.7% Minor 96.1% 94.9% 77.3% Severe 0.0% 0.0% 0.0% Structural Integrity Intact 57.3% 78.8% 92.2% Mildly broken 42.4% 21.2% 7.8% Broken 0.4% 0.0% 0.0% Semantic Coherence OK 62.0% 76.1% 92.2% Minor 38.0% 23.9% 7.8% Severe 0.0% 0.0% 0.0% Math State Preserved Yes 94.5% 97.3% 99.6% Borderline 5.5% 2.7% 0.4% No 0.0% 0.0% 0.0% SFT Suitability Yes 67.1% 85.9% 94.1% Borderline 32.9% 14.1% 5.9% No 0.0% 0.0% 0.0%
Table 3: Evaluation of pruned reasoning traces across multiple quality dimensions and keep fractions $\rho$.
E.4 Analysis
At $\rho=0.7$ (approximately 30% token reduction), degradation is primarily superficial. Most traces exhibit minor corruption, such as missing punctuation or dropped function words, but remain semantically coherent and structurally interpretable. Severe corruption is absent across all settings. Structural failures are rare (0.4% at $\rho=0.7$), and no examples exhibit a loss of mathematical state. In all cases, final answers remain present and numerically consistent. Suitability for SFT remains high even at aggressive pruning levels. At $\rho=0.7$, 67.1% of traces are directly usable, and the remaining ones are borderline, requiring only minor cleanup. No traces are deemed unusable. Manual inspection confirms that degradation is dominated by formatting artifacts rather than logical errors. Key reasoning steps, intermediate quantities, and final answers are preserved.
Takeaways:
These results suggest that greedy pruning preserves the semantic and mathematical integrity of reasoning traces even under substantial compression. While surface fluency degrades at lower keep fractions, this effect is largely orthogonal to the underlying reasoning and can be addressed with lightweight post-processing if desired. Overall, pruned traces remain suitable supervision for distillation and support the use of greedy pruning as a token-level compression method.
Prompt: Corruption Annotation SYSTEM PROMPT You are a strict annotator, evaluating the corruption of a pruned response compared to the original response. USER PROMPT TEMPLATE You are evaluating whether a pruned response is corrupted, using the following rubric. Your task has TWO STEPS, to be performed and output in order: Step 1 (Reasoning): Compare the PRUNED response to the ORIGINAL response step by step, in order. - For each step, directly compare the pruned and full response step (first with first, second with second, etc). - For each pair of steps, write a single line of analysis that considers: corruption (surface-level preservation, missing words, fragments, readability), structural integrity (step structure), semantic coherence (clarity of reasoning), math state (clarity/preservation of quantities), and suitability for SFT. - For each step, clearly state what is intact, what is mildly broken, and what is severely degraded according to the rubric. - Do not evaluate answer correctness; focus only on surface damage, structure, and coherence. - Write one line of analysis per step, in order, each on its own NEW LINE, precisely covering all rubric metrics for the step-by-step differences. Step 2 (Label): On a NEW LINE, output ONLY the final JSON (no extra text before or after it), using EXACTLY this schema and label space: { "corruption": "ok|minor|severe", "structural_integrity": "intact|mildly_broken|broken", "semantic_coherence": "ok|minor|severe", "math_state_preserved": "yes|borderline|no", "suitable_for_sft": "yes|borderline|no", "notes": "" } Rubric (for labeling): - corruption: * ok (readable/coherent), * minor (surface artifacts such as missing punctuation or truncated fragments but meaning recoverable), * severe (hard to parse or nonsensical) - structural_integrity: intact / mildly_broken / broken (whether step structure survives) - semantic_coherence: ok / minor / severe (whether the reasoning remains understandable to a human reader) - math_state_preserved: yes (main intermediate quantities/equations clearly present and consistent) / borderline (some degradation) / no (missing or inconsistent) - suitable_for_sft: yes (readable and structurally usable for SFT), borderline (usable but mildly degraded), no (too broken for practical SFT) - notes: If everything is perfect (corruption="ok" AND structural_integrity="intact" AND semantic_coherence="ok"), notes MUST be "". If suitable_for_sft="no", notes may still be "". Otherwise, use <= 15 words describing the surface-level defects. Remember: - Do NOT evaluate correctness of the answer. - Assess only readability, structure, and coherence of the PRUNED text versus the FULL text. Question: {{question}} Original Response (full): {{reason}} Pruned Response: {{pruned_reason}} First think step-by-step about each corresponding step in the pruned response and the original response. Then return the JSON. Return EXACTLY this JSON schema: { "corruption": "ok|minor|severe", "structural_integrity": "intact|mildly_broken|broken", "semantic_coherence": "ok|minor|severe", "math_state_preserved": "yes|no|unclear", "suitable_for_sft": "yes|borderline|no", "notes": "" }
Appendix F Student Behavior Under Pruned-Trace Distillation
In this section, we analyze how student models behave after training on greedily pruned reasoning traces. In particular, we study whether (i) students learn to generate shorter reasoning chains consistent with the pruned supervision, and (ii) whether this compression leads to degradation in structural integrity or semantic coherence.
F.1 Reasoning Length Adaptation
We first measure the length of reasoning traces generated by student models trained with different keep fractions $\rho\in\{1.0,0.9,0.8,0.7\}$. Table 4 reports the average number of generated reasoning tokens on the GSM8K validation set.
Keep Fraction $\rho$ Student Output Length
1.0 189 tokens
[…]
Table 4: Student reasoning lengths under different pruning levels. Values in parentheses indicate relative length normalized to the $\rho=1.0$ setting.
We observe that the student model closely matches the length distribution of the pruned training data. In particular, a keep fraction of $\rho=0.7$ yields approximately $30\%$ shorter reasoning traces at inference time. This indicates that the student does not revert to verbose reasoning, but instead internalizes the compressed reasoning style present in the supervision.
F.2 Structural and Semantic Properties of Student Outputs
We next evaluate whether compressed reasoning leads to degradation in output quality. Following the same protocol used for evaluating pruned training traces, we assess student-generated outputs along four dimensions: (i) structural integrity, (ii) semantic coherence, (iii) preservation of mathematical state, and (iv) overall suitability for SFT-style reasoning.
Across all pruning levels, we find that student outputs largely preserve the structure and semantics of the reasoning process. In particular, even at $\rho=0.7$, we do not observe cases where the reasoning becomes logically invalid or mathematically inconsistent. Instead, degradation is primarily stylistic, consisting of:
  • missing function words (e.g., articles, prepositions),
  • reduced punctuation,
  • occasional merging of reasoning steps.
These artifacts are consistent with those observed in the pruned training data and suggest that the student faithfully learns the distribution of compressed reasoning traces.
F.3 Interpretability of Compressed Reasoning
Despite the reduction in length, student-generated reasoning remains interpretable. In most cases, key intermediate quantities and computation steps are preserved, allowing the reasoning process to be followed with minimal effort. The primary effect of pruning is the removal of redundant narrative scaffolding rather than core computational content.
We emphasize that the goal of greedy pruning is not to produce stylistically optimal reasoning, but to preserve functional correctness under compression. If surface fluency is desired, this can be addressed orthogonally (e.g., via a lightweight rewriting or post-processing step), without altering the underlying token importance structure.
F.4 Summary
Overall, compressed-reasoning distillation yields student models that:
  • generate expected shorter reasoning traces,
  • preserve semantics and mathematical state,
  • exhibit only minor stylistic degradation.
These results suggest that greedy pruning provides a viable mechanism for training token-efficient reasoning models without materially compromising interpretability or correctness.